His recently departed Uncle Jimmy, may he rest in peace, has left Scott a 15-year annuity (15 annual payments). He will receive the first payment of $8,000 in 5 years from today, and payments will grow at a constant rate of 3% per year thereafter. If he is discounting at 6% (EAR), what is the present value of his inheritance?

Select one:

a. $73,911

b. $93,311

c. $69,727

d. $61,544

e. $127,530

f. $69,911

g. $98,909

h. No solution since r>g

NOTE: Please show all the work, without using EXCEL (Step by step with equations) Thanks!
Answer #1

Payment of $8000 will be received in 5 years from now. This mean the first cash flow will received at the end of 5th year.

So he will receive total 15 payments up to end of 19th year. Payments will grow at the rate 3% constantly every year.

The following table shows the cash flows in today terms:

Year

(1)

Receipt

(2)

PVF* @6%

(3)

Receipt*PVF*

(4) = (2)*(3)

1

0

0.943396 0
2 0 0.889996 0
3 0 0.839619 0
4 0 0.792094 0
5 8000 0.747258 5978.065
6 (8000*103%) = 8240 0.704961 5808.875
7 (8240*103%) = 8487.20 0.665057 5644.473
8 (8487.20*103%) = 8741.82 0.627412 5484.726
9 (8741.82*103%) = 9004.07 0.591898 5329.495
10 (9004.07*103%) = 9274.19 0.558395 5178.659
11 (9274.19*103%) = 9552.42 0.526788 5032.096
12 (9552.42*103%) = 9838.99 0.496969 4889.677
13 (9838.99*103%) = 10134.16 0.468839 4751.29
14 (10134.16*103%) = 10438.19 0.442301 4616.817
15 (10438.19*103%) = 10751.33 0.417265 4486.154
16 (10751.33*103%) = 11073.87 0.393646 4359.188
17 (11073.87*103%) = 11406.09 0.371364 4235.816
18 (11406.09*103%) = 11748.27 0.350344 4115.933
19 (11748.27*103%) = 12100.72 0.330513 3999.445
73910.71

* PVF = Present Value Factor

The Present Value of Inheritence is $73911. The answer is Option 'a'.